When two objects with masses of 200 kg and 500 kg were tested for gravitational attraction. Their centers were separated by a distance of 25 m. Find the force of gravitation between them. (ex: 6.57x10^10 N)


1
Expert's answer
2021-01-26T08:43:50-0500

The force of gravitation between them is


"F=G\frac{m_1m_2}{R^2},\\\space\\
F=6.673\cdot10^{-11}\frac{200\cdot500}{25^2}=1.07\cdot10^{-8}\text{ N}."